# GDGB: A Benchmark for Generative Dynamic Text-Attributed Graph Learning

### v1

@@ 4.2 Proposed Generative Metrics
To ensure a comprehensive assessment of graph structure, temporal dynamics, and textual quality of the generated DyTAG on TDGG and IDGG, we propose multifaceted evaluation protocols for DyTAG generation, integrating structural, textual, and graph embedding-based metrics. Detailed formulations and implementation specifics of the following metrics are provided in Section B.2.
Graph Structural Metric. We evaluate the structural validity of generated DyTAGs using two classical approaches: 1) Degree/Spectra MMD: We utilize the maximum mean discrepancy (MMD) with a radial basis function (RBF) kernel to measure the distribution distance between generated and ground-truth graphs (e.g., degree/spectral properties)MMD ; graphrnn ; EDGE . 2) Power-law Analysis: We assess power-law behavior in degree distributions using the Kolmogorov-Smirnov distance $D_{k}$ and the power-law exponent $\alpha$ powerlaw . Specifically, we assess the presence of power law validity by evaluating whether $D_{k}<0.15$ and $\alpha\in[2,3]$ for the generated graphs.
Textual Quality Metric. Inspired by the recent role-playing agent research RoleBench ; SocialBench ; charbox , we employ the LLM-as-Evaluator framework to assess the text quality in the generated DyTAGs. Specifically, the evaluation framework is conducted with five scoring criteria (Contextual Fidelity, Personality Depth, Dynamic Adaptability, Immersive Quality, Content Richness) tailored for DyTAG generation, scored on a 1–5 scale. Compared to embedding-based metrics used in prior DyTAG benchmarks (e.g., BERTScore bertscore in DTGB DTGB ), this evaluation framework offers two distinct advantages: 1) Unified multidimensional evaluation across diverse textual attributes, and 2) Preservation of semantic fidelity by avoiding information compression inherent in embedding representations.
Graph Embedding Metric. To enable a holistic and coupled evaluation of structural, temporal, and textual quality in DyTAG generation, we extend the JL-Metric jlmetric , designed for dynamic graph generation, to a better adaptation on DyTAGs by integrating textual node/edge features, resulting in a graph embedding-based indicator that quantifies the generation quality across three critical dimensions: topological patterns, temporal dynamics, and node/edge textual attributes. This metric condenses DyTAGs into a unified embedding space for the pairwise similarity computation between graph embeddings. The resulting scalar score quantifies global fidelity between generated and ground-truth DyTAGs by jointly measuring structural, temporal, and textual characteristics.
@@ 4.3 Generative Framework
Limitations of Previous Methods in DyTAG Generation. The novelty of our proposed TDGG and IDGG tasks presents unique challenges, as no existing method can directly address their requirements. For instance, the latest feature-supportive dynamic graph generation models like VRDAG VRDAG and DG-Gen DGGEN focus solely on generating node or edge representation features, but lack the capability to handle textual contents in DyTAGs. As far as we know, while GAG GAG is the only existing baseline that incorporates textual attributes, it is specifically designed for bipartite social graph generation and lacks generalizability to diverse DyTAG structures and domains. This gap motivates the development of a more universal framework tailored for DyTAG generation tasks.
GAG-General. Given the text-rich nature of DyTAGs, we propose GAG-General, an LLM-based multi-agent framework designed for DyTAG generation tasks, building upon the implementation of GAG GAG . Leveraging the text understanding and generation capabilities of LLMs, our framework extends the original GAG with three key enhancements: 1) Generalization: Support both bipartite and non-bipartite graph structures universally. 2) Multi-domain Compatibility: Abstract the DyTAG generation pipeline to enable seamless adaptation across diverse interaction scenarios without domain-specific customization. 3) Standardization: Establish standardized task formulations for TDGG and IDGG and incorporate holistic evaluation metrics, ensuring reproducible DyTAG generation.
Framework Details. Following GAG GAG , GAG-General employs carefully designed LLM-based agents for source and destination nodes, with each node agent maintaining a memory module to record historical neighbor interactions. This memory module effectively captures structural and temporal dynamics from the DyTAG by preserving contextual information about past interactions. We further incorporate an optional memory reflection mechanism, which leverages LLMs to distill the node memories into valuable summaries, akin to the message aggregation process in GNNs kipf2016semi ; peng2024beyond . For TDGG and IDGG, GAG-General implements an iterative, expansive DyTAG generation pipeline: in each iteration, the source node agent selects destination nodes based on its memory and contextual features. In the case of IDGG, new nodes are first generated and then updated to the node set before this selection step, ensuring the DyTAG’s inductive expansion. Subsequently, edges are generated between the selected node pairs. This iterative process continues until the final DyTAG is produced, seamlessly integrating node/edge generation and DyTAG evolution. See more details and pseudocode regarding the implementation of TDGG and IDGG based on GAG-General in Section B.1.
@@ 5.1 Experimental Settings
Baselines. 1) As our proposed GAG-General is an LLM-based multi-agent generative framework, for the DyTAG generation in TDGG and IDGG, we benchmark three open-source LLMs: DeepSeek-R1-Distill-Qwen-32B deepseek (referred to as DeepSeek), Llama-3-70B-Instruct llama3 (Llama), and Qwen2.5-72B-Instruct qwen2.5 (Qwen), as well as a closed-source model: GPT-4o-Mini gpt4omini (GPT). 2) For TDGG, which integrates traditional discriminative tasks from DGNNs in a generative paradigm, we compare against five state-of-the-art baselines, including JODIE JODIE , TGN TGN , CAWN CAWN , GraphMixer Graphmixer , and DyGFormer DyGLib . Performance is assessed on discriminative tasks such as node retrieval (link prediction) and edge classification to benchmark GAG-General against conventional DGNNs. 3) For IDGG, due to its novelty, the challenge of new node generation, and the text-rich nature of DyTAGs, we select two of the latest dynamic graph generation models that support node/edge feature modeling and generation: DG-Gen DGGEN and VRDAG VRDAG . See more details in Appendix C.
Implementation Details. 1) Both TDGG and IDGG share basic settings: the first 1,000 edges and the corresponding nodes serve as the initial seed DyTAG, and we set the edge generation size per round as 50. We employ the metrics introduced in Section 4.2 for evaluating the DyTAG generation quality on TDGG and IDGG, including graph structural metrics (e.g., Degree/Spectra MMD, power-law validity), textual quality scores, and a graph embedding-based indicator for joint structural-temporal-textual assessment. 2) For TDGG, we set the size of the target expanded DyTAG as 10,000 edges. To ensure fairness in discriminative task comparisons, we adopt a dataset split of 1,000/1,000/8,000 (train/val/test) for DGNNs based on the implementation of DyGLib DyGLib ; DTGB , and compare with the last 8,000 edges in the generated DyTAG from GAG-General. We use Hit@1 and Hit@10 for node retrieval (link prediction) tasks DTGB . For edge classification, we evaluate the models using weighted Precision, Recall, and F1 scores DTGB . 3) For IDGG, we set the size of the target expanded DyTAG as 2,000 edges, due to the additional stage of node generation. To ensure fair comparisons with dynamic graph generation baselines, we use the seed DyTAG with 1,000 edges as the input data for training DG-Gen and VRDAG and adjust the generation ratio in DG-Gen to generate 2,000 edges using its publicly available code DGGEN . As the current dynamic graph generation models’ lack of support for DyTAG’s textual features, our focus is on comparing the quality of generated graph structures and graph embeddings, rather than the texts. See more implementation details in Appendix C.
@@ 5.2 Transductive Dynamic Graph Generation
Results of TDGG. In TDGG, the structural quality of DyTAGs generated by GAG-General on GDGB datasets is shown in Table 2. Notably, the generated DyTAGs exhibit relatively small deviation from the ground-truths. For instance, the Degree/Spectra MMDs of most graphs in Table 2 are lower than 0.3, while six out of eight generated graphs satisfy the power-law validity criterion, demonstrating high-quality generation. To further investigate the impact of structural and textual information from historical neighbors on DyTAG generation, we conduct ablation studies among three configurations: w/o node memory, w/ node memory, and w/ node memory & reflection mechanism. As shown in Table 3 and 4, incorporating node memories or reflection mechanisms significantly improves the textual quality and the graph embedding-based indicator on most of the LLM backbones. This is attributed to the integration of structural and textual information via historical interactions (node memory) and the effective aggregation of node memories (reflection mechanism). Additionally, we conduct a scalability analysis of TDGG in Section D.2.
Table 2: The results on Degree MMD, Spectra MMD, $D_{k}$, $\alpha$, and power-law validity under TDGG with GPT as the LLM backbone. Full results on the other three LLM backbones are available in Table 13, 14, and 15.
@@
Table 3: The results on average textual quality scores under TDGG. M. and R. denote node memory and reflection mechanism, respectively. Full results on each scoring criterion are available in Table 16-24. The best and the runner-up scores are highlighted in bold and underlined fonts.
@@
Table 4: The results on the graph embedding metric under TDGG. See Table 25 for full results.
@@
Results of Discriminative Tasks in TDGG. For node retrieval, the results on Hit@1 and Hit@10 are presented in Table 5. For edge classification, the results are shown in Table 6. While existing DGNNs leverage complex modules to capture dynamic interaction patterns and achieve strong performance on discriminative tasks TGN ; Graphmixer ; DyGLib , their advantages diminish when trained solely on 1,000 edges. Notably, in Table 6, our generative framework outperforms DGNNs on most datasets, while the advantages on the node retrieval task are relatively less significant in Table 5. This highlights GAG-General’s superior ability to exploit the structural, temporal, and textual information in DyTAGs, whereas DGNNs exhibit limited capacity and over-reliance on large training input, resulting in poor generalization. These findings further validate our framework’s effectiveness in modeling DyTAGs, particularly in capturing complex dependencies for both discriminative and generative tasks.
Table 5: The results on Hit@1 and Hit@10 under the node retrieval task. Ours correspond to the best results of our proposed GAG-General among four LLM backbones. Full results on all LLM backbones and DGNN baselines are available in Table 26-29.
Hit@1 Hit@10
Sephora 0.285 0.045 0.023 0.065 0.028 0.386 0.259 0.206 0.309 0.184
Dianping 0.176 0.040 0.033 0.036 0.534 0.178 0.188 0.193 0.207 0.585
IMDB 0.142 0.015 0.009 0.005 0.008 0.159 0.123 0.110 0.105 0.111
WeiboTech 0.051 0.183 0.128 0.169 0.672 0.056 0.371 0.448 0.378 0.678
Table 6: The results on Precision, Recall, and F1 metric under the edge classification task. Ours correspond to the best results of our proposed GAG-General among four LLM backbones. Full results on all LLM backbones and DGNN baselines are available in Table 30- 32.
Datasets Models Ours JODIE TGN CAWN GraphMixer DyGFormer
Precision 0.798 0.463 0.452 0.442 0.443 0.471
Sephora Recall 0.790 0.661 0.672 0.661 0.684 0.687
F1 0.790 0.555 0.529 0.532 0.546 0.559
Precision 0.698 0.423 0.442 0.470 0.473 0.463
IMDB Recall 0.666 0.680 0.672 0.677 0.650 0.682
F1 0.672 0.531 0.542 0.551 0.546 0.533
5.3 Inductive Dynamic Graph Generation
Results of IDGG. In IDGG, the structural quality of DyTAGs generated by GAG-General on GDGB datasets is summarized in Table 7. Compared to TDGG, the generated DyTAGs in IDGG exhibit relatively greater deviation from the ground-truths since the new node generation increases task complexity. For instance, the Degree/Spectra MMDs of most graphs in Table 2 are greater than 0.2. However, the generated DyTAGs still retain key properties of the original, as five out of eight generated graphs still satisfy the power-law validity criterion, demonstrating reasonable structural fidelity. Similar to the TDGG, ablation studies confirm the importance of DyTAG’s structural and textual features. As shown in Table 8 and 9, node memories and reflection mechanisms significantly enhance textual quality and the graph embedding-based indicator, reaffirming the necessity of summarizing historical neighbor information for effective modeling of both graph structure and textual content. In addition, Table 10 presents a comparative analysis against existing dynamic graph generation models. The results reveal significant limitations in current approaches for modeling DyTAGs: in most cases, the generated graphs exhibit substantially lower quality in both structure fidelity and node/edge attribute richness than our generative framework. These findings underscore the urgent need for further advancements tailored to DyTAG generation. In addition, we perform a scalability analysis of IDGG in Section D.2.
Table 7: The results on Degree MMD, Spectra MMD, $D_{k}$, $\alpha$, and power-law validity under IDGG with GPT as the LLM backbone. Full results on the other three LLM backbones are available in Table 33, 34, and 35.
@@ 5.3 Inductive Dynamic Graph Generation
Table 8: The results on average textual quality scores under IDGG. Full results on each scoring criterion are available in Table 36-44.
@@
Table 9: The results on the graph embedding metric under IDGG. See Table 45 for full results.
@@
Table 10: The results on the graph structural and the graph embedding metrics under IDGG of our framework and current feature-supportive dynamic graph generation models. Ours correspond to the best results of our proposed GAG-General among four LLM backbones. See Table 46 for full results.
Sephora Dianping IMDB Cora
Ours VRDAG DG-Gen Ours VRDAG DG-Gen Ours VRDAG DG-Gen Ours VRDAG DG-Gen
Degree MMD 0.370 0.795 0.422 0.150 0.887 0.167 0.145 0.882 0.373 0.073 0.877 0.212
Spectra MMD 0.189 0.847 0.274 0.351 0.808 0.245 0.379 0.710 0.537 0.181 0.760 0.365
@@
Graph Embedding 0.661 0.011 0.228 0.769 0.024 0.517 0.729 0.095 0.293 0.828 0.053 0.056
Hub Node Analysis in IDGG. The generated DyTAGs in IDGG demonstrate a critical evolutionary pattern: while maintaining structural congruence with ground-truth graphs, they develop highly connected hub nodes with divergent textual attributes. Specifically, for generating a DyTAG with 5,000 edges on Sephora, the top three newly generated hub nodes in the generated DyTAG (with 74, 55, and 54 edges) include products like Vitamin E Moisturizer, Night Cream, and Facial Serum. In contrast, the ground-truth graph’s top three hub nodes (with 67, 66, and 27 edges) consist of products like Mineral Water Facial Spray, Exfoliating Treatment, and Oil Free Moisturizer. This divergence arises because IDGG mimics real-world graph evolution dynamics, ensuring structural fidelity while enabling the creation of new nodes with reasonable attributes that align with the underlying generative patterns of the real world. For instance, in recommendation systems, the hub nodes in the generated DyTAG can represent emerging products with high potential for virality, while those in the ground-truth graph correspond to established bestsellers. This capability positions DyTAG generation as a strategic tool for proactive decision-making in e-commerce and digital marketing. By identifying potential future hubs, platforms can prioritize resource allocation for product promotion, optimize advertising strategies, and anticipate market trends before they manifest in real-world data.

### v2

@@ 4.2 Proposed Generative Metrics
To ensure a comprehensive assessment of graph structure, temporal dynamics, and textual quality of the generated DyTAG on TDGG and IDGG, we propose multifaceted evaluation protocols for DyTAG generation, integrating structural, textual, and graph embedding-based metrics. Detailed formulations and implementation specifics of the following metrics are provided in Section˜D.3.
Graph Structural Metric. We evaluate the structural validity of generated DyTAGs using two classical approaches: 1) Degree/Spectra MMD: We utilize the maximum mean discrepancy (MMD) with a radial basis function (RBF) kernel to measure the distribution distance between generated and ground-truth graphs (e.g., degree/spectral properties)(Xiang et al., 2022; You et al., 2018; Chen et al., 2023). 2) Power-law Analysis: We assess power-law behavior in degree distributions using the Kolmogorov-Smirnov distance $D_{k}$ and the power-law exponent $\alpha$ (Clauset et al., 2009). Specifically, we assess the presence of power law validity by evaluating whether $D_{k}<0.15$ and $\alpha\in[2,3]$ for the generated graphs.
Textual Quality Metric. Inspired by the recent role-playing agent research (Wang et al., 2024b; Chen et al., 2024; Wang et al., 2024a), we employ the LLM-as-Evaluator framework to assess the text quality in the generated DyTAGs. Specifically, the evaluation framework is conducted with five scoring criteria (Contextual Fidelity, Personality Depth, Dynamic Adaptability, Immersive Quality, Content Richness) tailored for DyTAG generation, scored on a 1–5 scale. Compared to embedding-based metrics used in prior DyTAG benchmarks (e.g., BERTScore (Zhang et al., 2020) in DTGB (Zhang et al., 2024a)), this evaluation framework offers two distinct advantages: 1) Unified multidimensional evaluation across diverse textual attributes, and 2) Preservation of semantic fidelity by avoiding information compression inherent in embedding representations.
Graph Embedding Metric. To enable a holistic and coupled evaluation of structural, temporal, and textual quality in DyTAG generation, we extend the JL-Metric (Hosseini et al., 2025), designed for dynamic graph generation, to a better adaptation on DyTAGs by integrating textual node/edge features, resulting in a graph embedding-based indicator that quantifies the generation quality across three critical dimensions: topological patterns, temporal dynamics, and node/edge textual attributes. This metric condenses DyTAGs into a unified embedding space for the pairwise similarity computation between graph embeddings. The resulting scalar score quantifies global fidelity between generated and ground-truth DyTAGs by jointly measuring structural, temporal, and textual characteristics.
@@ 4.3 Generative Framework
Limitations of Previous Methods in DyTAG Generation. The novelty of our proposed TDGG and IDGG tasks presents unique challenges, as no existing method can directly address their requirements. For instance, the latest feature-supportive dynamic graph generation models like VRDAG (Li et al., 2024) and DG-Gen (Hosseini et al., 2024) focus solely on generating node or edge representation features, but lack the capability to handle textual contents in DyTAGs. As far as we know, while GAG (Ji et al., 2025) is the only existing baseline that incorporates textual attributes, it is specifically designed for bipartite social graph generation and lacks generalizability to diverse DyTAG structures and domains. This gap motivates the development of a more universal framework tailored for DyTAG generation tasks.
GAG-General. Given the text-rich nature of DyTAGs, we propose GAG-General, an LLM-based multi-agent framework designed for DyTAG generation tasks, building upon the implementation of GAG (Ji et al., 2025). Leveraging the text understanding and generation capabilities of LLMs, our framework extends the original GAG with three key enhancements: 1) Generalization: Support both bipartite and non-bipartite graph structures universally. 2) Multi-domain Compatibility: Abstract the DyTAG generation pipeline to enable seamless adaptation across diverse interaction scenarios without domain-specific customization. 3) Standardization: Establish standardized task formulations for TDGG and IDGG and incorporate holistic evaluation metrics, ensuring reproducible DyTAG generation. See more details of GAG-General in Section˜D.1.
Framework Details. Following GAG (Ji et al., 2025), GAG-General employs carefully designed LLM-based agents for source and destination nodes, with each node agent maintaining a memory module to record historical neighbor interactions. This memory module effectively captures structural and temporal dynamics from the DyTAG by preserving contextual information about past interactions. We further incorporate an optional memory reflection mechanism, which leverages LLMs to distill the node memories into valuable summaries, akin to the message aggregation process in GNNs (Kipf and Welling, 2016; Peng et al., 2024). For TDGG and IDGG, GAG-General implements an iterative, expansive DyTAG generation pipeline: in each iteration, the source node agent selects destination nodes based on its memory and contextual features. In the case of IDGG, new nodes are first generated and then updated to the node set before this selection step, ensuring the DyTAG’s inductive expansion. Subsequently, edges are generated between the selected node pairs. This iterative process continues until the final DyTAG is produced, seamlessly integrating node/edge generation and DyTAG evolution. See more details and pseudocode regarding the implementation of TDGG and IDGG based on GAG-General in Section˜D.2.
@@ 5.1 Experimental Settings
Baselines. 1) As our proposed GAG-General is an LLM-based multi-agent generative framework, for the DyTAG generation in TDGG and IDGG, we benchmark three open-source LLMs: DeepSeek-R1-Distill-Qwen-32B (DeepSeek-AI, 2025) (referred to as DeepSeek), Llama-3-70B-Instruct (AI@Meta, 2024) (Llama), and Qwen2.5-72B-Instruct (Team, 2024) (Qwen), as well as a closed-source model: GPT-4o-Mini (OpenAI, 2024) (GPT). 2) For TDGG, which integrates traditional discriminative tasks from DGNNs in a generative paradigm, we compare against five state-of-the-art baselines, including JODIE (Kumar et al., 2019), TGN (Rossi et al., 2020), CAWN (Wang et al., 2021a), GraphMixer (Cong et al., 2023), and DyGFormer (Yu et al., 2023). Performance is assessed on discriminative tasks such as node retrieval (link prediction) and edge classification to benchmark GAG-General against conventional DGNNs. 3) For IDGG, due to its novelty, the challenge of new node generation, and the text-rich nature of DyTAGs, we select two of the latest dynamic graph generation models that support node/edge feature modeling and generation: DG-Gen (Hosseini et al., 2024), VRDAG (Li et al., 2024), and TIGGER-I (Gupta et al., 2022). See more details in Appendix˜E.
Implementation Details. Both TDGG and IDGG share basic settings: the first 1,000 edges and the corresponding nodes serve as the initial seed DyTAG, and we set the edge generation size per round as 50. We employ the metrics introduced in Section˜4.2 for evaluating the DyTAG generation quality on TDGG and IDGG, including graph structural metrics, textual quality scores, and a graph embedding-based indicator for joint structural-temporal-textual assessment. See more implementation details in Appendix˜E.
@@ 5.2 Transductive Dynamic Graph Generation
Results of TDGG. In TDGG, GAG-General generates DyTAGs on GDGB with high structural fidelity, as shown in Table˜2. The generated graphs exhibit small deviations from ground-truths—most Degree/Spectra MMDs are below 0.3, and six of eight satisfy the power-law validity criterion, indicating high-quality generation. To assess the role of structural and textual information from historical neighbors, we ablate three configurations: w/o node memory, w/ node memory, and w/ node memory & reflection mechanism. As shown in Tables˜3 and 4, both node memory and the reflection mechanism significantly improve textual quality and graph embedding-based metrics across most LLM backbones, due to effective integration and aggregation of historical interaction information. Moreover, we experimentally verify the direct usability of generated graphs from TDGG in downstream tasks compared to the original graphs; experimental details are provided in Section˜F.5. Scalability analysis for TDGG is provided in Section˜F.1.
Table 2: The results on Degree MMD, Spectra MMD, $D_{k}$, $\alpha$, and power-law validity under TDGG with GPT as the LLM backbone. Full results on the other three LLM backbones are available in Table˜18, 19, and 20.
@@
Table 3: The results on average textual quality scores under TDGG. M. and R. denote node memory and reflection mechanism, respectively. Full results on each scoring criterion are available in Table˜21-29. The LLM (GPT) used in evaluation is independent of the LLM backbone used by GAG-General itself. See Section˜E.4 for more details. The best and the runner-up scores are highlighted in bold and underlined fonts.
@@
Table 4: The results on the graph embedding metric under TDGG. See Table˜30 for full results.
@@
Results of Discriminative Tasks in TDGG. For node retrieval, results on Hit@1 and Hit@10 are reported in Table˜36; for edge classification, in Table˜39. Although existing DGNNs employ sophisticated modules to model dynamic interactions and achieve strong performance on discriminative tasks (Yu et al., 2023), their effectiveness diminishes when training on only 1,000 edges. Notably, as shown in Table˜39, GAG-General outperforms DGNNs on most datasets, while gains on node retrieval are less pronounced (Table˜36). This demonstrates GAG-General’s superior ability to leverage structural, temporal, and textual information in DyTAGs, whereas DGNNs exhibit limited generalization due to their reliance on large-scale training data. These results further confirm GAG-General’s effectiveness in modeling DyTAGs, particularly in capturing dependencies across both discriminative and generative tasks.
Table 5: The results on Degree MMD, Spectra MMD, $D_{k}$, $\alpha$, and power-law validity under IDGG with GPT as the LLM backbone. Full results on the other three LLM backbones are available in Table˜40, 41, and 42.
@@
5.3 Inductive Dynamic Graph Generation
Results of IDGG. In IDGG, structural quality results are summarized in Table˜5. Generated DyTAGs show greater deviation from ground-truths than in TDGG—due to the added complexity of new node generation—yet retain key structural properties. For example, most Degree/Spectra MMDs exceed 0.2, but five of eight graphs still satisfy the power-law criterion, indicating reasonable fidelity. Similar to TDGG, ablation studies (Tables˜6 and 7) confirm that node memory and reflection enhance textual and structural quality, underscoring the importance of summarizing historical neighbor information. Moreover, Table˜8 compares our framework with existing dynamic graph generation models, revealing their significant limitations: generated graphs exhibit notably lower structural fidelity and poorer node/edge attribute richness. These results highlight the need for dedicated DyTAG generation methods. Additionally, we demonstrate that generated graphs from IDGG can serve as data augmentation to enhance model performance on inductive new nodes; see details in Section˜F.12. Scalability analysis for IDGG is also included in Section˜F.1.
Hub Node Analysis in IDGG. The generated DyTAGs in IDGG demonstrate a critical evolutionary pattern: while maintaining structural congruence with ground-truth graphs, they develop highly connected hub nodes with divergent textual attributes. Specifically, for generating a DyTAG with 2,000 edges on Sephora, the top three newly generated hub nodes in the generated DyTAG include products like Vitamin C Serum, Eye Cream, and Neck Cream. In contrast, the ground-truth graph’s top three hub nodes consist of products like Acne Control Clarifying Cleanser, Rosebud Perfume, and Moisturizing Lotion. The visualization of the hub node structures in the ground-truth and generated graphs is shown in Figures˜10 and 11. This divergence arises because IDGG mimics real-world graph evolution dynamics, ensuring structural fidelity while enabling the creation of new nodes with reasonable attributes that align with the underlying generative patterns of the real world. For instance, in recommendation systems, the hub nodes in the generated DyTAG can represent emerging products with high potential for virality, while those in the ground-truth graph correspond to established bestsellers. This capability positions DyTAG generation as a strategic tool for proactive decision-making in e-commerce and digital marketing. By identifying potential future hubs, platforms can prioritize resource allocation for product promotion, optimize advertising strategies, and anticipate market trends before they manifest in real-world data.
Table 6: The results on average textual quality scores under IDGG. Full results are available in Table˜43-51.
@@ 5.3 Inductive Dynamic Graph Generation
Table 7: The results on the graph embedding metric under IDGG. See Table˜52 for full results.
@@
Table 8: The results on the graph structural and the graph embedding metrics under IDGG of our framework and current feature-supportive dynamic graph generation models. Ours correspond to the best results of our proposed GAG-General among four LLM backbones. See Table˜53 for full results.
Sephora Dianping Cora
Ours VRDAG DG-Gen TIGGER-I Ours VRDAG DG-Gen TIGGER-I Ours VRDAG DG-Gen TIGGER-I
Degree MMD $\downarrow$ 0.370 0.795 0.422 0.622 0.150 0.887 0.167 0.446 0.073 0.877 0.212 0.372
Spectra MMD $\downarrow$ 0.189 0.847 0.274 0.687 0.351 0.808 0.245 0.341 0.181 0.760 0.365 0.389
@@
Graph Embedding $\uparrow$ 0.661 0.011 0.228 0.085 0.769 0.024 0.517 0.580 0.828 0.053 0.056 0.197